Find the pH and percent ionization of each HF solution (Ka for HF is 6.8 X 10^-4)
Please show your work, and explain. Thank you!
Co < CHE180 Review Exercise 16.77 Find the pH and percent ionization of each HF solution (Ka for HF is 6.8 x 104) PartA Find the pH of a 0.240 M solution. Express your answer to two decimal places. pH Submit Previous Answers Request Answer X Incorrect; One attempt remaining: Try Agairn Part B Find the percent dissociation of a 0.240 M HF solution. Express your answer using two significant figures
media%2Fdbc%2Fdbcaf0fd-8096-4f70-b27d-40
media%2F0e2%2F0e20be23-c0ed-4b81-ad67-f0
Answer #1


pH of weak acid is given by,

pH = 1/2(pKa - logC)

pH = 1/2(-log(6.8*10^-4) - log(0.24))

pH = 1.894

Percent of dissociation = square root of(Ka/C)

Percent of dissociation = square root of(6.8*10^-4/0.24) = 0.05322

pH of weak acid is given by,

pH = 1/2(pKa - logC)

pH = 1/2(-log(6.8*10^-4) - log(0.13))

pH = 2.027

Percent of dissociation = square root of(Ka/C)

Percent of dissociation = square root of(6.8*10^-4/0.13) = 0.0723

pH = 1/2(pKa - logC)

pH = 1/2(-log(6.8*10^-4) - log(4*10^-2))

pH = 2.283

Percent of dissociation = square root of(Ka/C)

Percent of dissociation = square root of(6.8*10^-4/4*10^-2) = 0.1303